OpenStudy (anonymous):

What is the value of (1/16) ^-3/2?

OpenStudy (saifoo.khan):

64

OpenStudy (anonymous):

How do you solve it?

OpenStudy (saifoo.khan):

Calc.

OpenStudy (anonymous):

Without a calc?

OpenStudy (saifoo.khan):

it can be solved.

OpenStudy (anonymous):

when you raise something to a negative value you flip the fraction so...

OpenStudy (saifoo.khan):

-ve exponent means invert the fraction..so 16^3/2 then , this can be written as. \[\Huge \sqrt{(16)^3}\] sqrt of 16 is 4. 4^3 4*4*4 = 64

OpenStudy (anonymous):

Ohh kk thanx.

OpenStudy (anonymous):

you have (16)^(3/2) and we know that squaring something is raising it to the "(1/2)" and we have that so we take the 3 out of the fraction and and save it for later. now all we have left is 16^(1/2)=sqrt(16)=4 now at the end we bring back the ^3 and cube the 4 thus getting 64